We have 4 hippos and we weigh them in pairs. After
weighing all possible pairs, we obtain the following list of
the weights of all pairs. Here is the list:
{356,362,374,374,386,392}
What is the sum of the weights of all the 4 hippos?

Answer by ankor@dixie-net.com(22740)

We have 4 hippos and we weigh them in pairs.
After weighing all possible pairs, we obtain the following list of
the weights of all pairs.
Here is the list:
{356, 362, 374, 374, 386, 392}
4 hippos a,b,c,d
;
a+b = 356
a+c = 362
a+d = 374
b+c = 374
b+d = 386
c+d = 392
:
Using the 1st two equations
a + 0 + c = 362
a + b + 0 = 356
-----------------subtraction eliminates a
-b + c = 6
c = (b+6)
using the 4th equation
b + c = 374
replace c with (b+6)
b + (b+6) = 374
2b = 374-6
2b = 368
b = 368/2
b = 184 kg is b's weight
then using the 1st equation
a + 184 = 356
a = 356 - 184
a = 172 kg is a's weight
:
Find c
c = 184 + 6
c = 190 kg is c's weight
:
find d
d + 190 = 392
d = 392 - 190
d = 202 kg is d's weight
"
find the sum
a = 172
b = 184
c = 190
d = 202
--------
sum 748 kg